OpenStudy (anonymous):

150/r = 600/r+150 Please help!

OpenStudy (anonymous):

I'm having issues with LCD etc... I keep getting the wrong answer!

OpenStudy (anonymous):

does r mean radius?

OpenStudy (anonymous):

no its rate

OpenStudy (anonymous):

ohhh no clue then sorry

OpenStudy (anonymous):

\[150/r=600/(r+150)\]\[(r)(150/r)=[600/(r+150)](r)\]\[150=600r/(r+150)\]\[150(r+150)=600r\]\[150r+22,500=600r\]\[22,500=600r-150r\]\[22,500=450r\]\[50=r\]